OpenStudy (kayla2001):

write a function rule for finding the amount of daily pay, p, in the following situation: a bus driver gets paid $100 each day plus $0.20 per kilometer, k.

OpenStudy (kayla2001):

@Indogot need help plzzzz

OpenStudy (anonymous):

im not a math guy. I'm history

OpenStudy (kayla2001):

o okay sorry to bother then

OpenStudy (anonymous):

its of @kayla2001

OpenStudy (kayla2001):

@ganeshie8 helppppppppppp

OpenStudy (kayla2001):

@neko_of_hearts I neeeed hlp please

OpenStudy (anonymous):

@Joel_the_boss i would help, but I'm not fully awake

OpenStudy (kayla2001):

ok thank for finding someone for me im both of yals fans

JoelTheBoss (joel_the_boss):

t=100+0.20x t is the total at the end of the day. x is the number of kilometer. Hope this helps. :)

OpenStudy (anonymous):

Mhm, welcome.

OpenStudy (kayla2001):

wait don't u mean p= and not t

OpenStudy (kayla2001):

that's not one of my answer choices

OpenStudy (anonymous):

He used different letter, oh that was easy....I could have..sorry @Joel_the_boss XD

OpenStudy (kayla2001):

wait that still not one of my choices

JoelTheBoss (joel_the_boss):

choices?

OpenStudy (kayla2001):

a.) 100=0.20 x k b.)p=0.20x 100x k c.) p =0.20k +100 d.)0.20k=100+p

OpenStudy (anonymous):

Well, which one do you think it is?

OpenStudy (kayla2001):

a

JoelTheBoss (joel_the_boss):

c is the answer lol 100 + another 0.20 cents for every kilometer and that would be the total.

OpenStudy (kayla2001):

..............................o

OpenStudy (anonymous):

Shhh I know XD don't give out answers Joel gosh!

OpenStudy (kayla2001):

lol

JoelTheBoss (joel_the_boss):

v_v I already explained and they said I was wrong so I gave them the same answer I already had but I explained it xD

OpenStudy (kayla2001):

thank yall some much medals anyone?

OpenStudy (kayla2001):

soooo should close it out?